# Chain-of-Thought in Neural Code Generation: From and For Lightweight Language Models

### v1

@@ 5.3 RQ3: Can $\ell$LMs benefit from the generated CoT?
To evaluate the effectiveness of generated CoTs in improving the code generation performance of $\ell$LMs, we first compare different models in the Self-CoT case (i.e., the CoTs generated by the model itself through the few-shot method in RQ1). We then consider the performance of $\ell$LMs (i.e., CodeGen, StarCoder and CodeT5+ as per Section 4.2) when using CoTs generated by CodeLlama, ChatGLM 130B, COTTON and the Teacher model. In particular, we compare the CoTs generated by CodeLlama and ChatGLM 130B with those generated by COTTON to ascertain if the CoTs generated by COTTON are more useful for $\ell$LMs. Additionally, we compare the gap between the CoTs generated by COTTON and those generated by the Teacher model. This comparison will provide insights into how closely COTTON can be to the more sophisticated Teacher model in generating CoTs. We use the Pass@1 metric to evaluate the effectiveness of using CoTs in improving the performance of $\ell$LMs, which will help us understand whether $\ell$LMs can effectively benefit from the instructions provided by CoT.
@@
The results are presented in Table VI, with the best result bolded and the second-best result underscored. Based on these results, we have the following three findings. First, all $\ell$LMs can effectively benefit from the instructions provided by CoTs to generate higher-quality code, assuming that the quality of the CoTs is guaranteed. This highlights the potential of using CoT in improving the performance of $\ell$LMs. Second, the performance of some models (such as CodeGen 2B and StarCoder 1B) may decrease under the guidance of CoTs generated by the self-CoT method. This finding supports the conclusion of RQ1, i.e., $\ell$LMs cannot independently generate high-quality CoTs. Third, among all the CoT generation methods, the CoTs generated by the Teacher Model show the most significant performance improvement for code generation by $\ell$LMs, followed by COTTON. This finding suggests that while COTTON cannot outperform the Teacher Model in most cases, it outperforms the other methods in terms of improving the code generation performance of $\ell$LMs, even outperforms LLMs like ChatGLM 130B.
@@
In our work, essentially we fine tune a base model CodeLlama to obtain COTTON. A legitimate question is why we choose CodeLlama, or whether other base models could yield better results. To further investigate this question, we take different base $\ell$LM models (i.e., GraphCodeBERT, CodeGPT-adapter, NatGen, CodeGeeX2) to tune them for CoT generation purposes. We then conduct experiments using the generated CoTs, where the results are presented in Table VII, with the best result in boldface and the second-best result underscored.
@@
Based on the results, it is observed that using CodeLlama as the base model leads to greater performance gains in code generation for the majority of $\ell$LMs, which justifies our choice and highlights the importance of choosing an appropriate base model when generating CoTs to enhance the code generation performance of $\ell$LMs. By selecting the most effective base model, COTTON can optimize the performance and effectiveness of $\ell$LMs in generating high-quality code.
$\ell$LMs can effectively benefit from the instructions provided by CoTs. It is observed that the code generation performance by $\ell$LMs can be improved with higher-quality CoTs. This highlights the potential of using CoTs in enhancing the performance of $\ell$LMs in the code generation task.
@@ 6.2 The impact of different prompt methods
Table VIII shows the performance of COTTON under different prompt methods on the HumanEval-CoT and OpenEval-CoT datasets, measured by BLEU-4 and Meteor scores. Among the prompt methods, Prefix Tuning consistently achieves the highest scores on both datasets. It demonstrates the effectiveness of adding task-specific prefixes to guide the model’s responses. Prompt Tuning and Alpaca Prompt also show competitive performance, indicating the benefits of using learned soft prompts or specific prompt templates.
@@
TABLE VIII: The performance of COTTON under different prompt methods
@@ 6.3 The impact of different hyper-parameter setting
TABLE IX: The performance of COTTON under different hyper-parameter settings
@@
The results are shown in Table IX, where we compare COTTON with other different hyper-parameter settings, and the performance is evaluated in terms of the BLEU-4 and METEOR metrics for both HumanEval-CoT and OpenEval-CoT. Based on the results, we find that a higher value of $Alpha$ tends to yield better performance. On the other hand, the impact of $R$ on COTTON performance is relatively less significant: varying its value does not lead to substantial performance improvement in terms of these evaluation metrics.
@@ 6.4 Threats to Validity
Threats to Construct Validity. The main construct threat is related to the metrics used in our automated evaluation. By treating CoT generation as a text generation problem, we utilize metrics based on term overlap (such as BLEU, METEOR, and ROUGE-L), which have been commonly used in similar studies on programming language processing [66, 67]. Moreover, we introduce the Consistency metric to assess alignment based on the nature of the code generation task under investigation. Furthermore, to ensure the generalizability of COTTON, we collect a new code generation dataset OpenEval. We manually design five test cases for each problem to ensure as much quality as possible for OpenEval. Finally, we employ Pass@1 and CoT-Pass@1 to evaluate the performance of code generation models. To complement automated evaluation, we also conducted a human study to further validate the effectiveness of our proposed approach. To guarantee the quality of our human study, we follow the human study methodology used in previous studies of similar software engineering tasks [68, 56].
@@ 7.2 Chain of Thought Generation
The methods have limitations in relying on LLMs with more than 100 billion parameters. Researchers have developed smaller language models via knowledge distillation. Ho et al. [95] introduced Fine-tune-CoT, which leverages GPT3(175B) as a reasoning teacher to enable complex reasoning in smaller models, thereby significantly reducing the model size requirements. Li et al. [96] proposed Symbolic Chain-of-Thought Distillation (SCoTD), a method for training smaller student models using rationalizations sampled from a much larger teacher model. This approach distills the reasoning capabilities of the larger model into smaller models. Shridhar et al. [97] utilized the step-by-step Chain-of-Thought (CoT) reasoning capabilities of larger models and distills these abilities into smaller models.
Our primary objective is to generate high-quality CoTs for code generation with manageable cost. Apart from the domain-specific features in constructing CoT, we design a stand-alone model of relatively small sizes dedicated to CoT generation which can improve the performance code generation.
@@ 8 Conclusion
Potential future research includes further improving the performance of COTTON. To this end, we plan to explore various potentially promising techniques, such as retrieval augmented generation, adversarial training, and contrastive learning. More generally, we believe it is important and promising to explore the lightweight language models in software engineering.
@@ Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (NSFC, No. 62372232), the Fundamental Research Funds for the Central Universities (No. NG2023005), the Collaborative Innovation Center of Novel Software Technology and Industrialization, and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. KYCX23_0396). T. Chen is partially supported by an oversea grant from the State Key Laboratory of Novel Software Technology, Nanjing University (KFKT2022A03, KFKT2023A04).

### v2

@@ 5.3 RQ3: Can $\ell$LMs benefit from the generated CoT?
5.3.1 Effect of self-generated CoTs on $\ell$LMs.
To assess the impact of generated CoTs on improving the code generation performance of LLMs (such as CodeGen, StarCoder, and CodeT5+ as discussed in Section 4.2), we initially compare different models in the Self-CoT scenario (i.e., the CoTs generated by the model itself using the few-shot method in RQ1). Furthermore, since Self-planning [13] and SCoT [14] have demonstrated that well-designed self-CoT prompts can effectively improve the performance of large models in code generation, we also include Think step-by-step [18], Self-planning [13], and SCoT [14] as baselines in our experiments.
@@
We use the Pass@1 metric to evaluate the effectiveness of using CoTs in improving the performance of $\ell$LMs, which can help understand whether $\ell$LMs can effectively benefit from the instructions provided by CoT. The results are presented in Table VII, with the best result highlighted in boldface and the second-best result underscored. We can observe that, under the guidance of the CoT generated by the Self-CoT method, the performance improvement of all $\ell$LMs is very limited, and even the performance of some $\ell$LMs may decrease. This finding supports the conclusion of RQ1, i.e., $\ell$LMs cannot independently generate high-quality CoTs.
5.3.2 Effect of CoTs generated by different LMs on $\ell$LMs.
We analyze the performance of $\ell$LMs when utilizing CoTs generated by CodeLlama, ChatGLM 130B, Gemini, and the Teacher model. Specifically, we compare the utility of CoTs generated by CodeLlama and the selected LLMs with those generated by COTTON to determine whether COTTON-generated CoTs are more beneficial for $\ell$LMs. The results are presented in Table VIII, with the best result highlighted in boldface and the second-best result underscored.
Our findings indicate that all $\ell$LMs can effectively leverage the guidance provided by CoTs to enhance the quality of generated code, assuming the CoTs’ quality is ensured. This emphasizes the potential of utilizing CoTs to improve the performance of $\ell$LMs.
Furthermore, we examine the disparity between CoTs generated by COTTON and those generated by the Teacher model. This comparison sheds light on how closely COTTON can approximate the more advanced Teacher model in CoT generation. The experimental results reveal that COTTON outperforms ChatGLM 130B and closely rivals Gemini and GPT-3.5-turbo in the CoT generation task for code generation.
5.3.3 Effect of CoTs generated by different base models on $\ell$LMs.
@@
In our work, essentially we fine-tune a base model CodeLlama to obtain COTTON. A legitimate question is why we choose CodeLlama, or whether other base models could yield better results. To further investigate this question, we take different base $\ell$LM models to tune them for CoT generation purposes. We then conduct experiments using the generated CoTs, where the results are presented in Table IX, with the best result in boldface and the second-best result underscored.
@@
Based on the results, it is observed that using CodeLlama as the base model leads to more significant performance improvements in code generation for the majority of $\ell$LMs, which justifies our choice and highlights the importance of choosing an appropriate base model when generating CoTs to enhance the code generation performance of $\ell$LMs. By selecting the most effective base model, COTTON can optimize the performance and effectiveness of $\ell$LMs in generating high-quality code.
Summary of RQ3 $\ell$LMs can effectively leverage the guidance provided by CoTs. This emphasizes the potential of utilizing CoTs to enhance the performance of $\ell$LMs in code generation tasks.
@@ 6 Discussions
6.1 Evaluating COTTON on LLMs
To further evaluate the capabilities of COTTON, we conduct experiments to assess its impact on LLMs, specifically focusing on gpt-3.5-turbo. Our goal is to investigate whether COTTON can enhance the performance of LLMs and to what extent this enhancement is achieved.
Gpt-3.5-turbo serves as a representative model in code generation and is one of the most state-of-the-art LLMs available. We consider its performance for code generation in zero-shot scenarios and then under the guidance of various COTs, including the CoTs generated by COTTON, the CoTs self-generated by Self-planning [13], SCoT [14], and the Self-CoT method proposed in our study.
TABLE X: Evaluating COTTON and other CoT prompt methods on gpt-3.5-turbo
Corpus Pass@1 CoT-Pass@1
COTTON Self-Planning SCoT Self-CoT
HE 56.10 74.39 77.44 76.83 77.44
OE 26.97 43.26 42.70 44.38 45.51
HE-p 57.32 76.22 78.66 78.66 79.27
Table X shows the impact of COTTON and other self-generated CoT methods on gpt-3.5-turbo. The experimental results show that the gpt-3.5-turbo model demonstrates a significant improvement in code generation performance when guided by various CoTs. When comparing the zero-shot scenarios with the guided code generation, we observe a substantial increase in Pass@1, indicating that the incorporation of CoTs, including those generated by COTTON, positively influences the model’s ability to generate more accurate code. Notably, it even exceeds the performance of the GPT4 zero-shot scenario on the HumanEval dataset (67.0). Moreover, the Self-CoT method proposed in our study can perform better than Self-Planning and SCoT on all datasets.
TABLE XI: The performance comparison of zero-shot, fine-tuning and COTTON with StarCoder series models.
StarCoder Pass@1 CoT-Pass@1
Zero-shot Fine-tune Instruction-tune COTTON COTTON w. Fine-tune COTTON w. Instruction-tune
HE 1B 12.80 14.63 ($\uparrow$ 14.30%) 16.46 ($\uparrow$ 28.59%) 25.00 ($\uparrow$ 95.31%) 25.00 ($\uparrow$ 95.31%) 28.05 ($\uparrow$ 119.14%)
3B 17.07 22.56 ($\uparrow$ 32.16%) 24.36 ($\uparrow$ 42.71%) 30.49 ($\uparrow$ 78.62%) 31.71 ($\uparrow$ 85.76%) 35.37 ($\uparrow$ 107.21%)
7B 21.95 25.61 ($\uparrow$ 16.67%) 26.83 ($\uparrow$ 22.23%) 37.20 ($\uparrow$ 69.48%) 38.41 ($\uparrow$ 74.99%) 40.24 ($\uparrow$ 83.33%)
16B 34.10 37.80 ($\uparrow$ 10.85%) 39.02 ($\uparrow$ 14.43%) 43.90 ($\uparrow$ 28.74%) 44.51 ($\uparrow$ 30.53%) 46.95 ($\uparrow$ 37.68%)
OE 1B 8.99 8.43 11.24 ($\uparrow$ 25.03%) 17.42 ($\uparrow$ 93.77%) 17.98 ($\uparrow$ 100%) 20.22 ($\uparrow$ 124.92%)
3B 11.24 18.54 ($\uparrow$ 64.95%) 19.66 ($\uparrow$ 74.91%) 19.10 ($\uparrow$ 69.93%) 20.79 ($\uparrow$ 84.96%) 22.47 ($\uparrow$ 99.91%)
7B 23.03 24.16 ($\uparrow$ 4.91%) 26.40 ($\uparrow$ 14.63%) 33.15 ($\uparrow$ 43.94%) 33.71 ($\uparrow$ 46.37%) 35.96 ($\uparrow$ 56.14%)
16B 27.53 29.78 ($\uparrow$ 8.17%) 31.46 ($\uparrow$ 14.28%) 39.33 ($\uparrow$ 42.86%) 40.45 ($\uparrow$ 46.93%) 42.70 ($\uparrow$ 55.10%)
HE-p 1B 13.41 14.63 ($\uparrow$ 9.10%) 16.46 ($\uparrow$ 22.74%) 26.22 ($\uparrow$ 95.53%) 26.83 ($\uparrow$ 100.07%) 28.66 ($\uparrow$ 113.72%)
3B 17.07 22.56 ($\uparrow$ 32.16%) 24.36 ($\uparrow$ 42.71%) 31.71 ($\uparrow$ 85.76%) 31.71 ($\uparrow$ 85.76%) 36.59 ($\uparrow$ 114.35%)
7B 22.56 26.22 ($\uparrow$ 16.22%) 26.83 ($\uparrow$ 22.23%) 38.41 ($\uparrow$ 70.26%) 38.41 ($\uparrow$ 74.99%) 40.24 ($\uparrow$ 83.33%)
16B 34.10 38.41 ($\uparrow$ 12.64%) 39.02 ($\uparrow$ 14.43%) 44.51 ($\uparrow$ 30.53%) 44.51 ($\uparrow$ 30.53%) 46.95 ($\uparrow$ 37.68%)
6.2 Ablation study
(a) Results in HumanEval-CoT
(b) Results in OpenEval-CoT
Figure 5: The performance of COTTON by whether using the Consistency Checker
The data collection methodology for CodeCoT-9k relies on heuristic rules and multi-agent alignment. The heuristic rules (including Code Filtering, Document Filtering, and Similarity Filtering) have been widely used in training various LLMs (such as DeepSeek-Coder [66], StarCoder [40], CodeT5+ [41], and WizardCoder [67]), and their importance has been demonstrated. The multi-agent alignment, particularly the Quality Checker, has been proven by Gunasekar et al. [33] to be beneficial for training models with educational value. However, the specific contribution of the Consistency Checker to the overall quality of the dataset remains unclear. To evaluate the impact of the Consistency Checker, we conduct an ablation experiment. When the Consistency Checker is not applied, the dataset consists of 13k examples. In this experiment, we train COTTON separately on CodeCoT-9k and a 13k dataset without Consistency Checker filtering.
The ablation results of the Consistency Checker are shown in Figure 5 where we can observe that the use of the Consistency Checker has a significant impact on the performance of COTTON. Specifically, in terms of the Consistency metric, without the Consistency Checker for filtering, COTTON shows a relative performance decrease of 5.23% on HumanEval-CoT and 4.69% on OpenEval-CoT.
6.3 Comparing COTTON with other tune methods
In this subsection, we first compare COTTON with traditional fine-tuning method using the StarCoder-series models as examples. The empirical study [36] shows that LoRA usually makes the most favorable trade-off between cost and performance in code generation tasks. Hence we mainly compare COTTON with LoRA in our discussion. During the traditional fine-tuning process, natural language is used as input, and code is generated as output.
Given that existing state-of-the-art instruction-tuning methods usually involve additional data, such as human feedback [68] and compiler feedback [69, 70], which can significantly improve the performance of downstream tasks of the model, we follow the method of Zheng et al. [71] and add CoT in the form of multi-turn dialogues during instruction-tuning to explore the impact of instruction-tuning on the performance of the model.
The results shown in Table XI demonstrate that the combination of StarCoder-7B with COTTON can exceed the performance of StarCoder-16B in zero-shot scenarios and can even achieve results comparable to a fine-tuned/instruction-tuned StarCoder-16B model. In addition, we find that multi-tune instruction-tuning method outperform traditional fine-tuning method, which is a finding similar to that of Zheng et al. [71]. We also find that COTTON, as an independent CoT generation model, can be combined with the fine-tuned/instruction-tuned model to further improve the performance of the model in code generation tasks.
Furthermore, when considering GPU across model deployment, inference, and training with float16 precision (as shown in Table XII), we carefully set parameters including a maximum input and output length of 256 and a batch size of 1. It is important to highlight that fine-tuning a 7B model with LoRA already pushes the limits on a single consumer-grade GPU (GeForce RTX series). For individual developers, fine-tuning a 16B model would require multiple professional-grade GPUs (Quadro series and Tesla series), leading to significantly higher hardware costs compared to fine-tuning a 7B model. In contrast, COTTON enables performance enhancement across multiple models without the need of fine-tuning individual model, which is highly desirable.
TABLE XII: The comparison of GPU memory usage with StarCoder series models.
StarCoder Deploy Inference Training (Lora)
1B 2.34 GB $\sim$3.50 GB $\sim$12.00 GB
3B 6.15 GB $\sim$7.27 GB $\sim$16.00 GB
7B 14.72 GB $\sim$15.90 GB $\sim$23.00 GB
16B 31.88 GB $\sim$33.00 GB $\sim$40.00 GB
@@ 6.5 The impact of different prompt methods
Table XIII shows the performance of COTTON under different prompt methods on the HumanEval-CoT and OpenEval-CoT datasets, measured by BLEU-4 and Meteor scores. Among the prompt methods, Prefix Tuning consistently achieves the highest scores on both datasets. It demonstrates the effectiveness of adding task-specific prefixes to guide the model’s responses. Prompt Tuning and Alpaca Prompt also show competitive performance, indicating the benefits of using learned soft prompts or specific prompt templates.
@@
This discussion shows that considering a variety of different prompt methods during model training can have unexpected benefits in improving model performance. Therefore, when targeting different downstream tasks, it is necessary to carefully try different prompt methods to maximize the potential of the model.
TABLE XIII: The performance of COTTON under different prompt methods
@@ 6.6 The impact of different hyper-parameter setting
TABLE XIV: The performance of COTTON under different hyper-parameter settings
@@
The results are shown in Table XIV, where we compare COTTON with other different hyper-parameter settings, and the performance is evaluated in terms of the BLEU-4 and METEOR metrics for both HumanEval-CoT and OpenEval-CoT. Based on the results, we find that a higher value of $Alpha$ tends to yield better performance. On the other hand, the impact of $R$ on COTTON performance is relatively less significant: varying its value does not lead to substantial performance improvement in terms of these evaluation metrics.
@@ 6.7 Threats to Validity
Threats to Construct Validity. The main construct threat is related to the metrics used in our automated evaluation. By treating CoT generation as a text generation problem, we utilize metrics based on term overlap (such as BLEU, METEOR, and ROUGE-L), which have been commonly used in similar studies on programming language processing [80, 81]. Moreover, we introduce the Consistency metric to assess alignment based on the nature of the code generation task under investigation. Furthermore, to ensure the generalizability of COTTON, we collect a new code generation dataset OpenEval. We manually design five test cases for each problem to ensure as much quality as possible for OpenEval. Finally, we employ Pass@1 and CoT-Pass@1 to evaluate the performance of code generation models. It is noticed that the Pass@k metric would significantly increase the computational cost of the evaluation, especially when k is large, in our study we decided to focus only on the Pass@1 metric in the evaluation, which reflects the actual ability of the model to generate the correct code for a given input in one go, which is the most critical aspect of functional correctness. To complement automated evaluation, we also conducted a human study to validate the effectiveness of our proposed approach further. To guarantee the quality of our human study, we follow the human study methodology used in previous studies of similar software engineering tasks [82, 61].
@@ 7.2 Chain of Thought Generation
The methods have limitations in relying on LLMs with more than 100 billion parameters. Researchers have developed smaller language models via knowledge distillation. Ho et al. [108] introduced Fine-tune-CoT, which leverages GPT3(175B) as a reasoning teacher to enable complex reasoning in smaller models, thereby significantly reducing the model size requirements. Li et al. [109] proposed Symbolic Chain-of-Thought Distillation (SCoTD), a method for training smaller student models using rationalizations sampled from a much larger teacher model. This approach distills the reasoning capabilities of the larger model into smaller models. Shridhar et al. [110] utilized the step-by-step Chain-of-Thought (CoT) reasoning capabilities of larger models and distilled these abilities into smaller models.
Our primary objective is to generate high-quality CoTs for code generation at a manageable cost. Apart from the domain-specific features in constructing CoT, we design a stand-alone model of relatively small sizes dedicated to CoT generation which can improve the performance code generation.
7.3 Multi-agent Collaboration
Multi-agent collaboration refers to a framework where multiple autonomous agents interact with each other in a shared environment. These agents can be program scripts, software bots, or robots, each with their capabilities, goals, and perceptions [111]. They can communicate, cooperate, compete, or negotiate with each other to achieve complex goals or solve problems. LLMs within multi-agent collaboration systems is an emerging area of research in the deep learning community [112].
For instance, Zhang et al. [113] proposed ProAgent, a system designed for robotic tasks that analyzes the current context, anticipates teammates’ intentions, and formulates strategies based on this reasoning. Chen et al. [114] developed VisualGPT, which leverages vision-based Pretrained Language Models for image captioning tasks. In terms of code generation, Huang et al. [115] proposed AgentCoder, a novel solution comprising a multi-agent framework with the programmer agent, the test designer agent, and the test executor agent.
Our proposed multi-agent alignment-based cleaning method contains Quality Checker, CoT Generator, and Consistency Checker. Quality Checker serves the purpose of evaluating and filtering the educational significance of the code, while the CoT Generator focuses on generating the corresponding CoTs. Additionally, the Consistency Checker plays a crucial role in assessing and filtering the semantic consistency of the generated CoTs with the code. These three agents possess unique capabilities and task goals, enabling them to communicate and collaborate effectively to produce high-quality CoTs.
@@ 8 Conclusion
Potential future research includes further improving the performance of COTTON. We plan to extend COTTON to other programming languages and explore various potentially promising techniques, such as retrieval augmented generation, adversarial training and contrastive learning. Moreover, based on the results of our study, it is important and promising to explore the lightweight language models for other software engineering tasks.
@@ Acknowledgements
The authors would like to thank the editors and the anonymous reviewers for their insightful comments and suggestions, which can substantially improve the quality of this work. This work was partially supported by the National Natural Science Foundation of China (NSFC, No. 62372232), the Fundamental Research Funds for the Central Universities (No. NG2023005), the Collaborative Innovation Center of Novel Software Technology and Industrialization, the Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. KYCX23_0396), and the Short-term Visiting Program of Nanjing University of Aeronautics and Astronautics for Ph.D. Students Abroad (No. 240501DF16). T. Chen is partially supported by oversea grants from the State Key Laboratory of Novel Software Technology, Nanjing University (KFKT2022A03, KFKT2023A04).